OpenStudy (anonymous):

What is the dy/dx of x^5= -xy^3?

OpenStudy (anonymous):

product rule and chain rule for this one

OpenStudy (anonymous):

yeah mate i still don't undestand. i'm a bit lost

OpenStudy (anonymous):

start with \[5x^4=-(y^3+3xy^2y')\] and solve for \(y'\)

OpenStudy (anonymous):

you are thinking that \(y=f(x)\) even though you have not solved for it visualize it as \[x^5=-xf^3(x)\]

OpenStudy (anonymous):

I am coming up with -(6x^4 + y^3)/3xy^2

OpenStudy (anonymous):

\[5x^4=-(y^3+3xy^2y')\] is the first step or \[5x^4=-y^3-3xy^2y'\]

OpenStudy (anonymous):

i think you just have a typo there with the \(6\)

OpenStudy (anonymous):

otherwise your answer is right

OpenStudy (anonymous):

-(5x^4 + y^3)/3xy^2 right?

OpenStudy (anonymous):

looks good to me